# TQml Simulator: Optimized Simulation of Quantum Machine Learning

### v2

@@ 1 Introduction
Here, we aim to develop an optimized quantum simulator specifically for QML applications. For each gate layer, we benchmarked a set of applicable approaches and found that the optimal simulation method for a particular layer often depends on the number of qubits in the circuit. Based on this observation, we propose an approach to build an ”Optimized” simulator of a QML circuit that exploits the optimal simulation techniques chosen for each gate layer individually in order to minimize the forward or combined forward and backward execution times for the available hardware, such as multithreaded CPUs, GPUs, or TPUs. Our approach plays a role similar to that of modern software compilers that target a specific hardware architecture, but it acts one layer higher: instead of rearranging low-level instructions as, e.g, hardware-centric quantum-simulation compilers do [18], it chooses the most efficient algorithmic recipe for each gate layer for given computational resources. We named the obtained simulator as Terra Quantum Machine Learning (TQml) Simulator. By implementing TQml Simulator using a PyTorch back-end, we show that it outperforms PennyLane default.qubit simulator by up to a factor of 10, depending on the circuit, the number of qubits, the batch size of the input data, and the hardware used. The remainder of the paper is structured as follows. In Sec. 2, we present the simulation techniques for various layers of gates and provide numerical benchmark results comparing these techniques. In Sec. 3, we describe our method for creating the optimized TQml Simulator with PyTorch back-end for QML models and benchmark its performance against the PennyLane default.qubit simulator using a representative circuit, evaluating both input data parallelization and full training iteration (forward and backward passes). Then, in Sec. 4 we also benchmark our TQml simulator with JAX back-end and analyze its performance in comparison with PyTorch back-end and other simulators. Finally, in Sec. 5, we summarize our findings and conclude our work.
@@ 2.1 Methods
Z $\boldsymbol{\vee}$cdot $\boldsymbol{\vee}$cdot 1
CZ $\boldsymbol{\vee}$cdot $\boldsymbol{\vee}$cdot 2
Rz $\boldsymbol{\vee}$cdot $\cdot \boldsymbol{\vee}$ 1
Rzz $\boldsymbol{\vee}$cdot $\cdot \boldsymbol{\vee}$ 2
GPI $\boldsymbol{\vee}$cdot $\cdot$ $\cdot$ 1
S $\boldsymbol{\vee}$cdot $\cdot$ $\cdot$ 1
CNOT $\cdot \boldsymbol{\vee}$${\vee}$cdot 2
X $\cdot \boldsymbol{\vee}$${\vee}$cdot 1
H $\cdot \cdot \boldsymbol{\vee}$cdot 1
Ry $\cdot \cdot \boldsymbol{\vee}$${\vee}$ 1
Rx $\cdot \cdot \cdot \boldsymbol{\vee}$ 1
Rot $\cdot \cdot \cdot \boldsymbol{\vee}$ 1
GPI2 $\cdot \cdot \cdot \boldsymbol{\vee}$ 1
MS $\cdot \cdot \cdot \boldsymbol{\vee}$ 2
@@ 2.2.2 Sequential Einsum
Hardware-efficient circuit ansätze commonly used in QML applications [20] rely on local quantum operations—such as single-qubit rotations and two-qubit entangling gates—that are applied uniformly across all qubits. In this context, applying such local gates using Einstein summation can be preferred over other approaches.
@@ 2.4.1 Eigenphase Computation
where $K_{Rz}$ is a $2^{n}\times n$ matrix with entries of ±$\pm 1$. Each row of $K_{Rz}$ represents a unique binary configuration of the $n$ qubits, with $+1$ corresponding to a 00 and $-1$ corresponding to a $1$. To generate $K_{Rz}$, one starts with a $2^{n}\times n$ binary counting matrix $J$ and applies the elementwise transformation:
@@ 2.4.2 Diagonal Tensor Product
where the superscript indicates the corresponding qubit. Computing this tensor product requires $42^{n+1}-$ 4 multiplications, which is $\mathcal{O}(2^{n})$. Once the full diagonal is computed, applying the gate layer to the state vector reduces to an elementwise multiplication:
@@ 2.5 H-Rz Expansion
[…]
@@ Appendix B Hardware-specific Simulators: IBM and IonQ
Native gates on IBM’s Eagle-type superconducting processors include ECR, X, Sx, and $R_{z}$. Using these gates, we constructed a QDI circuit layer [21] (see also the main text) as shown in Fig. 11(a). The benchmark results presented in Fig. 12 demonstrate that the TQml Simulator achieves a speedup of up to an order of magnitude for circuits with up to approximately 15 qubits, depending on the batch size and the hardware used for simulation. However, as the number of qubits increases, the performance advantage of our simulator converges towards that of the default.qubit simulator.

### v3

@@ 1 Introduction
Here, we aim to develop an optimized quantum simulator specifically for QML applications [18, 19, 20]. For each gate layer, we benchmarked a set of applicable approaches and found that the optimal simulation method for a particular layer often depends on the number of qubits in the circuit. Based on this observation, we propose an approach to build an ”Optimized” simulator of a QML circuit that exploits the optimal simulation techniques chosen for each gate layer individually in order to minimize the forward or combined forward and backward execution times for the available hardware, such as multithreaded CPUs, GPUs, or TPUs. Our approach plays a role similar to that of modern software compilers that target a specific hardware architecture, but it acts one layer higher: instead of rearranging low-level instructions as, e.g, hardware-centric quantum-simulation compilers do [21], it chooses the most efficient algorithmic recipe for each gate layer for given computational resources. We named the obtained simulator as Terra Quantum Machine Learning (TQml) Simulator. By implementing TQml Simulator using a PyTorch back-end, we show that it outperforms PennyLane default.qubit simulator by up to a factor of 10, depending on the circuit, the number of qubits, the batch size of the input data, and the hardware used. The remainder of the paper is structured as follows. In Sec. 2, we present the simulation techniques for various layers of gates and provide numerical benchmark results comparing these techniques. In Sec. 3, we describe our method for creating the optimized TQml Simulator with PyTorch back-end for QML models and benchmark its performance against the PennyLane default.qubit simulator using a representative circuit, evaluating both input data parallelization and full training iteration (forward and backward passes). Then, in Sec. 4 we also benchmark our TQml simulator with JAX back-end and analyze its performance in comparison with PyTorch back-end and other simulators. Finally, in Sec. 5, we summarize our findings and conclude our work.
@@ 2.1 Methods
[…]
@@ 2.2.2 Sequential Einsum
Hardware-efficient circuit ansätze commonly used in QML applications [23] rely on local quantum operations, such as single-qubit rotations and two-qubit entangling gates, that are applied uniformly across all qubits. In this context, applying such local gates using Einstein summation can be preferred over other approaches.
@@ 2.4.1 Eigenphase Computation
where $K_{Rz}$ is a $2^{n}\times n$ matrix with entries of $\pm 1$. Each row of $K_{Rz}$ represents a unique binary configuration of the $n$ qubits, with +1+1 corresponding to a 0 and 1-1 corresponding to a $1$. To generate $K_{Rz}$, one starts with a $2^{n}\times n$ binary counting matrix $J$ and applies the elementwise transformation:
@@ 2.4.2 Diagonal Tensor Product
where the superscript indicates the corresponding qubit. Computing this tensor product requires $2^{n+1}-4$ multiplications, which is $\mathcal{O}(2^{n})$. Once the full diagonal is computed, applying the gate layer to the state vector reduces to an elementwise multiplication:
@@ 2.5 H-Rz Expansion
$\mathrm{GPI2}(\theta)$ $=R_{z}(-\theta)\,R_{x}(\pi/2)\,R_{z}(\theta).$ (4)
In simulating these decompositions, we employed the optimal techniques for a given number of qubits as indicated in Fig. 2 for H layers and in Fig. 4 for Rz layers. Similar to $H$ gate, $R_{x}(\pi/2)$ in the $\mathrm{GPI2}(\theta)$ decomposition is constant for a given number of qubits, and we used the same optimal techniques except the Real Unitary Operation since $R_{x}(\pi/2)$ is complex. As shown in Fig. 5, this H-Rz Expansion method surprisingly outperforms other approaches, such as Einsum, for up to 10 qubits in the case of $R_{y}$ and $R_{x}$ gates, and it happens to be optimal for a layer of general rotation Rot and GPI2 gates.
@@ Appendix B Hardware-specific Simulators: IBM and IonQ
Native gates on IBM’s Eagle-type superconducting processors include ECR, X, Sx, and $R_{z}$. Using these gates, we constructed a QDI circuit layer [24, 20] (see also the main text) as shown in Fig. 11(a). The benchmark results presented in Fig. 12 demonstrate that the TQml Simulator achieves a speedup of up to an order of magnitude for circuits with up to approximately 15 qubits, depending on the batch size and the hardware used for simulation. However, as the number of qubits increases, the performance advantage of our simulator converges towards that of the default.qubit simulator.